# CleverDistiller: Simple and Spatially Consistent Cross-modal Distillation

### v1

@@ 4.1 Experiments setup
Though most cross-modal image-to-LiDAR distillation works claim to pre-train for only 50 epochs, some recent works such as SuperFlow [83] utilize 3 LiDAR sweeps per sample, increasing the effective length and compute cost of the training by a factor of three (effectively 150 epochs). We characterize this using effective epochs in Tab. 4. For fairer comparison, we show results of our method for both 50 and 100 epochs of pre-training. CleverDistiller50 provides a direct comparison to previous methods that used only one LiDAR sweep per sample while CleverDistiller100 serves as a more efficient comparison to Superflow. After pre-training we fine-tune the models for 100 epochs.
@@
PPKT [51] arXiv’21 ViT-B 50 39.95 40.91 53.21 60.87 66.22 74.07 44.09 47.57
@@
PPKT [51] arXiv’21 ViT-L 50 41.57 42.05 55.75 61.26 66.88 74.33 45.87 47.82
@@ 4.3 Performance Comparison
Leveraging data representation learning can significantly reduce the dependency on data annotations. We evaluate CleverDistiller against prior SOTA methods on three prominent datasets: nuScenes, SemanticKITTI, and Waymo, under scenarios with limited annotations for few-shot fine-tuning. As illustrated in Tab. 4 and Tab. 5, CleverDistiller demonstrates SOTA performance across all fine-tuning tasks as well as on linear probing on almost every scenario on both MinkUnet [19] and PTV3 [75]. Especially when fine-tuning on low data amounts on nuScenes we observe strong improvements of up to 10% mIoU, e.g. for 1% nuScenes data and ViT-L DINOv2 teacher. This performance indicates strong generalization of our learned feature representations, making our work highly relevant for use cases with low data availability. The tables also show that models pretrained through representation learning consistently surpass those initialized randomly, highlighting the benefits of data pretraining. Additionally, distillations from larger 2D networks yield consistent enhancements which is consistent with observations in previous works.
@@
LP PPKT50 [51] MinkU-34 183.44 78.15 30.65 35.42 28.12 29.21 32.82 19.52 28.01 20.71 28.06
@@ 4.4 Domain generalization
We test our method, pre-trained on nuScenes, on the ScriKITTI, Rellis-3D, SemSTF, and DAPS-3D datasets to show its domain generalization capabilities, as shown in Tab. 6. Each dataset presents unique challenges, and our method consistently performs well. Our approach consistently achieved the highest scores in most cases compared to other SOTA pretraining methods. For instance, on the ScriKITTI dataset, our method reached 40.77 mIoU with 1% data and 53.31 with 10% data, outperforming all other methods. Similarly, on the Rellis-3D dataset, our method achieved 56.23 mIoU with 1% data and 56.60 with 10% data, again leading the performance charts. Even on the SemSTF and DAPS-3D datasets, our method demonstrated strong results, particularly excelling with 82.58 mIoU at 50% data and 86.15 at 100% data on DAPS-3D. These results underscore the robustness and generalization of our approach in 3D semantic segmentation tasks.
@@ 4.5 Robustness
The robustness of 3D perception models against unprecedented conditions directly correlates with the model’s applicability to real-world applications. We compare our method with prior models in the nuScenes-C dataset from the Robo3D [40] benchmark and show results in Tab. 7. Our method demonstrates superior performance, particularly in corrupted scenarios where other methods struggle, such as when the LiDAR data is much sparser.

### v2

@@ 4.1 Experiments setup
Though most cross-modal image-to-LiDAR distillation works claim to pre-train for only 50 epochs, some recent works such as SuperFlow [83] utilize 3 LiDAR sweeps per sample, increasing the effective length and compute cost of the training by a factor of three (effectively 150 epochs). We characterize this using effective epochs in Sec. 4.1. For fairer comparison, we show results of our method for both 50 and 100 epochs of pre-training. CleverDistiller50 provides a direct comparison to previous methods that used only one LiDAR sweep per sample while CleverDistiller100 serves as a more efficient comparison to Superflow. After pre-training we fine-tune the models for 100 epochs.
@@
black      lightgray PPKT [51] arXiv’21 ViT-B 50 39.95 40.91 53.21 60.87 66.22 74.07 44.09 47.57
@@
black      lightgray PPKT [51] arXiv’21 ViT-L 50 41.57 42.05 55.75 61.26 66.88 74.33 45.87 47.82
@@ 4.3 Performance Comparison
Leveraging data representation learning can significantly reduce the dependency on data annotations. We evaluate CleverDistiller against prior SOTA methods on three prominent datasets: nuScenes, SemanticKITTI, and Waymo, under scenarios with limited annotations for few-shot fine-tuning. As illustrated in Sec. 4.1 and Tab. 5, CleverDistiller demonstrates SOTA performance across all fine-tuning tasks as well as on linear probing on almost every scenario on both MinkUnet [19] and PTV3 [75]. Especially when fine-tuning on low data amounts on nuScenes we observe strong improvements of up to 10% mIoU, e.g. for 1% nuScenes data and ViT-L DINOv2 teacher. This performance indicates strong generalization of our learned feature representations, making our work highly relevant for use cases with low data availability. The tables also show that models pretrained through representation learning consistently surpass those initialized randomly, highlighting the benefits of data pretraining. Additionally, distillations from larger 2D networks yield consistent enhancements which is consistent with observations in previous works.
@@
black    lightgray LP PPKT50 [51] MinkU-34 183.44 78.15 30.65 35.42 28.12 29.21 32.82 19.52 28.01 20.71 28.06
@@ 4.4 Domain generalization
We test our method, pre-trained on nuScenes, on the ScriKITTI, Rellis-3D, SemSTF, and DAPS-3D datasets to show its domain generalization capabilities, as shown in Sec. 4.3. Each dataset presents unique challenges, and our method consistently performs well. Our approach consistently achieved the highest scores in most cases compared to other SOTA pretraining methods. For instance, on the ScriKITTI dataset, our method reached 40.77 mIoU with 1% data and 53.31 with 10% data, outperforming all other methods. Similarly, on the Rellis-3D dataset, our method achieved 56.23 mIoU with 1% data and 56.60 with 10% data, again leading the performance charts. Even on the SemSTF and DAPS-3D datasets, our method demonstrated strong results, particularly excelling with 82.58 mIoU at 50% data and 86.15 at 100% data on DAPS-3D. These results underscore the robustness and generalization of our approach in 3D semantic segmentation tasks.
@@ 4.5 Robustness
The robustness of 3D perception models against unprecedented conditions directly correlates with the model’s applicability to real-world applications. We compare our method with prior models in the nuScenes-C dataset from the Robo3D [40] benchmark and show results in Sec. 4.3. Our method demonstrates superior performance, particularly in corrupted scenarios where other methods struggle, such as when the LiDAR data is much sparser.